Circular No. 8944

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2008 J5 (GARRADD) G. J. Garradd reports his discovery of a comet with a 40′′ coma and a

broad 20′′ tail toward the southwest on CCD survey images obtained with the 0.5-m Uppsala Schmidt telescope (discovery observation tabulated below). Following posting on the Minor Planet Center’s ‘NEOCP’ webpage, C. Jacques and E. Pimentel (Belo Horizonte, Minas Gerais, Brazil, 0.30-m f/3 reflector) write that their CCD images from May 4.3 UT show a condensed coma 30′′ in diameter. J. C. Pelle found a diffuse coma of diameter 20′′ with a bright central condensation and a faint tail 20′′ long in p.a. 220^{o} on images taken on May 14.54–14.57 by N. Teamo with a 0.41-m f/8 reflector at Punaauia, French Polynesia. R. Ligustri (Talmassons, Udine, Italy) found a 24′′ coma on stacked CCD exposures obtained remotely with a 0.25-m f/6 reflector at Moorook, South Australia, on May 14.74–14.76.

2008 UT  α_{2000}  δ_{2000}  Mag.
May 13.73380  22^{h}28^{m}50.^{s}70  −42^{o}14′48′′.1  16.7

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2008-J76.

T = 2008 May 3.175 TT  ω = 329.^{o}123
Ω = 285.207 }2000.0
q = 2.01712 AU  i = 96.824

COMET P/2008 J2 (BESHORE) Additional observations have shown this comet (cf. IAUC 8941) to be

of short period; the following orbital elements are from MPEC 2008-J68:

T = 2008 Apr. 24.018 TT  ω = 143.^{o}165
e = 0.27988  Ω = 97.432 }2000.0
q = 2.45485 AU  i = 10.622

a = 3.40895 AU n^{o} = 0.156593 P = 6.29 years Visual total-magnitude and coma-diameter estimates: May 7.43 UT,

14.1, 0′.3 (A. Hale, Cloudcroft, NM, 0.41-m reflector); 9.97, 13.0, ∼ 1′ (M. L. Paradowski, Dabrowa, Poland, 0.20-m reflector); 12.07, 12.7, 1′.2 (J. J. Gonzalez, Leon, Spain, 0.20-m reflector).

2008 May 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green